OpenStudy (anonymous):

Use a half angle formula to solve for the identity below: (cos(2x))^2

OpenStudy (anonymous):

\[\cos^2\alpha=\frac{1}{2}\left(1+\cos2\alpha\right)\] which gives you \[\cos2\alpha=2\cos^2\alpha-1\] Substitute \(\alpha=2x\).

OpenStudy (anonymous):

what if it was (cos(6x))^4

OpenStudy (anonymous):

\[\cos^46x=\left(\cos^26x\right)^2\] You'd apply the half angle formula to rewrite \(\cos^26x=\cdots\), then you'd have \(\left(\cdots\right)^2\). The inside would be a binomial, which could then be expanded.